Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can make as many corrections as she wants. Create the formula for the sum of this geometric series and explain your steps in solving for the maximum grade Diana can receive. Identify this as converging or diverging.
Answer by richwmiller(17219)

She lost 35 points
She can get 3/10 back
35*3/10=105/10=10.5 points maximum for new grade of 75.5 maximum.
converging on 75.5 or 10.5 depending on your viewpoint
65+.3*n=s